A 197-kg crate is pushed horizontally with a force of 702 N. If the coefficient of friction is 0.20, calculate the acceleration of the crate.
1
Expert's answer
2013-11-25T08:51:14-0500
Let us find friction force:
F_f = m*g*mu = 197*9.8*0.2 = 386 N

Differences of forces will cause acceleration:

ma = F-F_f
a = (F-F_f)/m = (702-386)/197 = 1.6 m/s^2